OpenStudy (anonymous):

The equation: (x^2 + 4)(x^2 - 4) What I got: X^16-16x^4. Now on my assignment, I have to fill in the blanks. I did so, and I had two extra blanks. I am supposed to fill in all of them. Help?

OpenStudy (anonymous):

What method is being used in here?

OpenStudy (cwrw238):

(x^2 + 4)(x^2 - 4) multiply second bracket by x^2 then by + 4 = x^4 - 4x^2 + 4x^2 - 16 = x^4 - 16 the middle terms cancel out

OpenStudy (anonymous):

Oh ok, thank you

OpenStudy (cwrw238):

here we use the distribute law

OpenStudy (cwrw238):

yw